OpenStudy (anonymous):

Add or subtract -4x/x+7 - 8/x+7 PLEASE HELP.

OpenStudy (anonymous):

you have to do -4x-8 and keep the denominator

OpenStudy (anonymous):

subtract in the numerator \[\frac{-4x}{x+7}-\frac{8}{x+7}=\frac{-4x-8}{x+7}\]

OpenStudy (anonymous):

Thank you!